OpenStudy (anonymous):

Find the inverse of the function. f(x) = cube root x/8 -4

hartnn (hartnn):

is that \(f(x)=(\sqrt[3]x/8) -4\) ?

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{\frac{ x }{ 8 }}-4\]

hartnn (hartnn):

oh, ok to get the inverse let y= f(x) and interchange 'x' and 'y'

OpenStudy (anonymous):

just solve for x and at the end swap x and y.

hartnn (hartnn):

\(y= \sqrt[3]{\dfrac{x}{8}}-4\\x= \sqrt[3]{\dfrac{y}{8}}-4\) solve for 'y' ! can you ?

OpenStudy (anonymous):